what is the 10th # in this sequence 1,2,4: there are 5 possible answers please help
Found 2 solutions by Alan3354, scott8148:
Answer by Alan3354(69443)

The sequence might be 1,2,4,8,16,32... successive terms times 2
Or 1,2,4,7,11,16,22... the difference between successive terms increasing by 1
Or 1,2,4,5,7,8 ... the difference alternating between 1 and 2
Or other sequences.
Not sufficient data.

Answer by scott8148(6628)

the value of successive terms is doubled

the value of any term is ___ 2^(n-1) ___ where n is the number of the term (1st, 2nd, 3rd, etc.)

the value of the 10th term is ___ 2^(10-1) = 2^9 = 512